# Diff-Instruct*: Towards Human-Preferred One-step Text-to-image Generative Models

### v2

@@ Abstract
In this paper, we introduce the Diff-Instruct* (DI*), an image data-free approach for building one-step text-to-image generative models that align with human preference while maintaining the ability to generate highly realistic images. We frame human preference alignment as online reinforcement learning using human feedback (RLHF), where the goal is to maximize the reward function while regularizing the generator distribution to remain close to a reference diffusion process. Unlike traditional RLHF approaches, which rely on the KL divergence for regularization, we introduce a novel score-based divergence regularization, which leads to significantly better performances. Although the direct calculation of this preference alignment objective remains intractable, we demonstrate that we can efficiently compute its gradient by deriving an equivalent yet tractable loss function. Remarkably, we used Diff-Instruct* to train a Stable Diffusion-XL-based 1-step model, the 2.6B DI*-SDXL-1step text-to-image model, which can generate images of a resolution of $1024\times 1024$ with only 1 generation step. DI*-SDXL-1step model uses only 1.88% inference time and 29.30% GPU memory cost to outperform 12B FLUX-dev-50step significantly in PickScore, ImageReward, and CLIPScore on Parti prompt benchmark and HPSv2.1 on Human Preference Score benchmark, establishing a new state-of-the-art benchmark of human-preferred 1-step text-to-image generative models. Besides the strong quantitative performances, extensive qualitative comparisons also confirm the advantages of DI* in terms of maintaining diversity, improving image layouts, and enhancing aesthetic colors. We have released our industry-ready model on the homepage: https://github.com/pkulwj1994/diff_instruct_star.
@@ 1 Introductions
Generative models have made substantial progress in recent years, largely improving the content creation and editing across various domains (Karras et al., 2020; Nichol & Dhariwal, 2021; Poole et al., 2022; Kim et al., 2022; Tashiro et al., 2021; Meng et al., 2021; Couairon et al., 2022; Ramesh et al., 2022; Esser et al., 2024; Oord et al., 2016; Ho et al., 2022; Brooks et al., 2024; Zhang et al., 2023a; Xue et al., 2023; Luo & Zhang, 2024; Luo et al., 2023b; Zhang et al., 2023b; Feng et al., 2023; Deng et al., 2024; Luo et al., 2024d; Geng et al., 2024; Wang et al., 2024; Pokle et al., 2022).
@@
Particularly, two types of generative models have gained significant attention, diffusion models and one-step generators. Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020), or score-based generative models (Song et al., 2020), progressively corrupt data with diffusion processes and then train models to approximate the score functions of the noisy data distributions across varying noise levels. The learned score functions can generate high-quality samples by denoising the noisy samples through reversed stochastic differential equations. While DMs can produce good outputs, they often require a large number of model evaluations, which limits their efficiency in applications.
@@
Different from diffusion models, one-step generators (Luo et al., 2024c; a; Zheng & Yang, 2024; Kang et al., 2023; Sauer et al., 2023a; Yin et al., 2024; Zhou et al., 2024a) have emerged as a highly efficient alternative to multi-step diffusion models. Unlike DMs, one-step generative models directly map latent noises as inputs to generate data samples with a single neural network forward pass. This 1-step generation mechanism makes them ideal for real-time applications such as text-to-image and text-to-video generations on edge devices. Many existing works have demonstrated the leading performances of one-step text-to-image generators (Luo et al., 2024c; Zhou et al., 2024a; Yin et al., 2024) by employing diffusion distillation (Luo, 2023) as well as GAN techniques(Goodfellow et al., 2014; Brock et al., 2018; Karras et al., 2020; Sauer et al., 2023b). However, these works only focus on matching the 1-step model’s distributions with pre-trained diffusion models or ground truth data, without considering the critical challenge of teaching one-step text-to-image models to satisfy human preferences, which is one of the most important needs in the age of human-centric AI.
@@
Figure 1: Generated $1024\times 1024$ images from 1-step 2.6B DI*-SDXL model. The DI*-SDXL-1step model outperforms FLUX-dev-50step(flu, ) and SD3.5-large-28step(sd, 3) in human preference metrics on the Parti prompt benchmark and HPSv2.1 scores. Please zoom in to check the details.
@@
To resolve this issue, we introduce Diff-Instruct* (DI*), a novel approach that can train one-step text-to-image generators to have strong human-preference alignment while still maintaining the ability to generate diverse and realistic high-resolution images. We frame the human-preference alignment problem as maximizing the rewards with a score-based divergence constraint. This yields generated samples that not only have improved human reward but also adhere to user prompts. Different from traditional RLHF methods (Christiano et al., 2017; Ouyang et al., 2022; Luo, 2024), which rely on Kullback-Leibler (KL) divergence for distribution regularization, we find score-based regularization maintains better diversities, preventing an aligned 1-step model from stupidly generating painting-like images that have high human reward scores but poor diversity.
In our experiments, we use Diff-Instruct* to train 1-step text-to-image models based on SDXL(Podell et al., 2023) reference diffusion. We introduced the DI*-SDXL-1step model, an industry-ready model with 2.6B parameters and a UNet architecture. We evaluate our DI*-SDXL-1step model with other leading models extensively on both Parti prompts (Yu et al., 2022), COCO prompts, and Human Preference Score (HPS) v2.1 benchmarks. On Parti-prompt benchmark, the DI*-SDXL-1step model shows a very strong performance, outperforming 12B FLUX-dev-50step on PickScore, ImageReward, and CLIPScore, with only 1.88% inference time and 29.30% GPU memory cost. On COCO prompt datasets, the DI*-SDXL-1step goes on par with FLUX-dev and SD3.5-large diffusion models. On the Human Preference Score benchmarks, the 2.6B DI*-SDXL-1step model achieves a SoTA overall HPSv2.1 score of 31.19, demonstrating very strong human preference performances among all open-sourced models including Diffusio-DPO, DMD2, and FLUX-dev models. These solid results establish the Diff-Instruct* as a simple yet powerful approach for preference alignment of text-to-image models with SoTA open-sourced 1-step models.
We summarize our contribution in the following points:
@@
  • We introduced Diff-Instruct* a novel approach for preference alignment of 1-step generative models with theoretical guarantees based on a novel score-based PPO paradigm;
@@
  • We introduce two novel approaches to incorporate traditional classifier-free guidance into human preference alignment through explicit-implicit reward decoupling.
@@
  • With extensive evaluations on three standard benchmarks across different alignment configurations, our best open-sourced 2.6B DI*-SDXL-1step text-to-image model, a leading text-to-image model with a resolution of $1024\times 1024$ that significantly outperforms previous leading FLUX-dev-50step model with only 1.88% inference time.
@@ 2 Preliminary
Here we introduce preliminary knowledge and notations about diffusion models. Let $q_{0}(\bm{x})=q_{d}(\bm{x})$ be the data distribution. The goal of generative modeling is to train models to generate new samples $\bm{x}\sim q_{0}(\bm{x})$. Under mild conditions, the forward diffusion process of a diffusion model can transform initial distribution $q_{0}$ towards some simple noise distribution,
@@
$\mathcal{L}_{DSM}(\varphi)=\int_{t=0}^{T}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim q_{0},\atop\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\|\bm{s}_{\varphi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$ (2.2)
The weighting function $\lambda(t)$ controls the importance of the learning at different time levels and $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ denotes the conditional transition of the forward diffusion (2.1). After training, $\bm{s}_{\varphi}(\bm{x}_{t},t)\approx\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ is a good approximation of marginal scores of diffused data distribution.
3 Preference Alignment using Diff-Instruct*
@@
Our basic setting is that we have a known human reward function $r(\bm{x}_{0},\bm{c})$, which encodes the human preference for an image $\bm{x}_{0}$ and corresponding text description $\bm{c}$. Besides, we also have a pre-trained diffusion model which will later act as a reference distribution $p_{ref}(\bm{x}_{0})=q_{0}(\bm{x}_{0})$. The reference diffusion model is specified by the score function $\bm{s}_{q_{t}}(\bm{x}_{t})\coloneqq\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$, where $q_{t}(\bm{x}_{t})$’s is the underlying distribution diffused at time $t$ according to (2.1). The pre-trained diffusion model represents the ground-truth data distribution that is used to prevent the aligned 1-step model from seeking high human reward but losing the ability to generate realistic images.
@@
Our goal is to train a human-preferred one-step text-to-image model $g_{\theta}$ that generates images by directly mapping a random noise $\bm{z}\sim p_{z}$ to obtain $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$, conditioned on the input text $\bm{c}\sim\mathcal{C}$. We want the 1-step model’s output distribution, $p_{\theta}(\bm{x}_{0}|\bm{c})$, to maximize the expected human rewards while maintaining the ability to generate realistic images. We force $p_{\theta}(\bm{x}_{0}|\bm{c})$ to not diverge from $p_{ref}(\cdot)$ by using a divergence regularization term. Let $\bm{D}(\cdot,\cdot)$ be a distribution divergence. For any fixed prompt $\bm{c}$, the training objective is defined as:
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}|\bm{c})}\big{[}-\alpha r(\bm{x}_{0},\bm{c})\big{]}+\bm{D}(p_{\theta},p_{ref})$ (3.1)
@@
Here $\alpha$ is a coefficient that balances reward influences and $\bm{D}(\cdot)$ acts as a regularization term. When using Kullback-Leibler divergence as the regularization term, the objective (3.1) turns to the online Proximal Policy Gradient (PPO) algorithm(Schulman et al., 2017), which has great success in reinforcement learning from human feedback (RLHF) for large language models (Ouyang et al., 2022) and one-step text-to-image models alignment (Luo, 2024). Particularly, recently (Luo, 2024) has shown that traditional KL-PPO can fairly train human-preferred one-step text-to-image models. However, KL divergence is notorious for mode-seeking behavior, which makes Luo (2024) very easy to collapse to some distribution modes, such as painting-like images, that have very high human reward but lack diversity. To address this issue, we find that general score-based divergences have an appealing diversity-preserving property that shows great successes in 1-step diffusion distillation (Luo et al., 2024c; Zhou et al., 2024b). Based on such observations, we propose Diff-Instruct*, a novel practical algorithm based on an innovative Score-based online PPO paradigm that uses a general score-based divergence as regularization for reinforcement learning using human feedback.
Figure 2: Overall workflow of developing 1-step text-to-image generative models using Diff-Instruct*. The workflow includes stages of pre-training, reward modeling, and preference alignment.
3.1 General Score-based Divergences
@@
Inspired by score-implicit matching (Luo et al., 2024c) that revealed the advantages of score-based divergences over KL divergences, we define the regularization term $\bm{D}(p_{\theta},p_{ref})$ via the following general score-based divergence. Assume $\mathbf{d}:\mathbb{R}^{d}\to\mathbb{R}$ is a scalar-valued proper distance function (i.e., a non-negative function that satisfies $\forall\bm{x},\mathbf{d}(\bm{x})\geq 0$ and $\mathbf{d}(\bm{x})=0$ if and only if $\bm{x}=\bm{0}$). Given a parameter-independent sampling distribution $\pi_{t}$ that has large distribution support, we can formally define a time-integral score divergence as
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})\coloneqq\int_{t=0}^{T}w(t)\mathbb{E}_{\bm{x}_{t}\sim\pi_{t}}\bigg{\{}\mathbf{d}(\bm{s}_{p_{\theta,t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t}))\bigg{\}}\mathrm{d}t,$ (3.2)
@@
3.2 Diff-Instruct*
Recall that $g_{\theta}$ is a one-step model, therefore samples from $p_{\theta}$ can be implemented through a direct mapping $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$. With the score-based regularization term (3.2), for each given text prompt $\bm{c}$, we can formally write down our training objective to minimize as:
@@ 3.2 Diff-Instruct*
$+\int_{t=0}^{T}w(t)\mathbb{E}_{\bm{x}_{t}|\bm{x}_{0}\atop\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\bigg{\{}-\mathbf{d}^{\prime}(\bm{y}_{t})\bigg{\}}^{T}\bigg{\{}\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}\mathrm{d}t\bigg{]}$
@@
Under mild assumptions, if we take the sampling distribution in (3.2) as $\pi_{t}=p_{\operatorname{sg}[\theta],t}$, then the $\theta$ gradient of (3.3) is the same as the objective (3.4): $\frac{\partial}{\partial\theta}\mathcal{L}_{Orig}(\theta)=\frac{\partial}{\partial\theta}\mathcal{L}_{DI*}(\theta)$.
In practice, we can use another assistant diffusion model $\bm{s}_{\psi}(\bm{x}_{t},t)$ to approximate the 1-step model’s score function $\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})$ pointwise, which was also done in the literature of diffusion distillations works such as Luo et al. (2024b); Zhou et al. (2024a; b); Luo et al. (2024c); Yin et al. (2023; 2024). Therefore, we can alternate between 1) updating the assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t},t)$ using 1-step model generated samples (which are efficient) and 2) updating the 1-step model by minimizing the tractable objective (3.4). We name our training method that minimizes the objective $\mathcal{L}_{DI*}(\theta)$ in (3.4) the Diff-Instruct* because it is inspired by Diff-Instruct(Luo et al., 2024b) and Diff-Instruct++(Luo, 2024) that involves an additional diffusion model and a reward model to train one-step t2i models.
3.3 Two Ways of Incorporating Classifier-free Guidance
Input: prompt dataset $\mathcal{C}$, 1-step model $g_{\theta}(\bm{x}_{0}|\bm{z},\bm{c})$, prior distribution $p_{z}$, reward model $r(\bm{x},\bm{c})$, reward model scale $\alpha_{rew}$, CFG reward scale $\alpha_{cfg}$, reference diffusion model $\bm{s}_{ref}(\bm{x}_{t}|c,\bm{c})$, assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})$, forward diffusion $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ (2.1), assistant diffusion updates rounds $K_{TA}$, time distribution $\pi(t)$, diffusion model weighting $\lambda(t)$, student loss time weighting $w(t)$.

### v3

# David and Goliath: Small One-step Model Beats Large Diffusion with Score Post-training

@@ Abstract
We propose Diff-Instruct*(DI*), a data-efficient post-training approach for one-step text-to-image generative models to improve its human preferences without requiring image data. Our method frames alignment as online reinforcement learning from human feedback (RLHF), which optimizes the one-step model to maximize human reward functions while being regularized to be kept close to a reference diffusion process. Unlike traditional RLHF approaches, which rely on the Kullback-Leibler divergence as the regularization, we introduce a novel general score-based divergence regularization that substantially improves performance as well as post-training stability. Although the general score-based RLHF objective is intractable to optimize, we derive a strictly equivalent tractable loss function in theory that can efficiently compute its gradient for optimizations. We introduce DI*-SDXL-1step, which is a 2.6B one-step text-to-image model at a resolution of $1024\times 1024$, post-trained from DMD2 w.r.t SDXL. Our 2.6B DI*-SDXL-1step model outperforms the 50-step 12B FLUX-dev model in ImageReward, PickScore, and CLIP score on the Parti prompts benchmark while using only 1.88% of the inference time. This result clearly shows that with proper post-training, the small one-step model is capable of beating huge multi-step diffusion models. Our model is open-sourced at this link: https://github.com/pkulwj1994/diff_instruct_star. We hope our findings can contribute to human-centric machine learning techniques.
Machine Learning, ICML
@@ 1 Introductions
Generative models have made substantial progress in recent years, largely improving the creative content creation across various domains (Karras et al., 2020; Nichol & Dhariwal, 2021; Poole et al., 2022; Kim et al., 2022; Tashiro et al., 2021; Meng et al., 2021; Couairon et al., 2022; Ramesh et al., 2022; Esser et al., 2024; Oord et al., 2016; Ho et al., 2022; Brooks et al., 2024; Zhang et al., 2023a; Xue et al., 2023; Luo & Zhang, 2024; Luo et al., 2023b; Pokle et al., 2022; Zhang et al., 2023b; Feng et al., 2023; Deng et al., 2024; Luo et al., 2024d; Geng et al., 2024; Wang et al., 2024).
@@
Among these advancements, two types of generative models have gained significant attention, diffusion models (DMs) and one-step generators. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), or score-based generative models (Song et al., 2020), progressively corrupt data with diffusion processes and then train models to approximate the score functions of the noisy data distributions across varying noise levels. The learned score functions can generate high-quality samples by denoising the noisy samples through reversed stochastic differential equations. While DMs can produce good outputs, they often require a large number of model evaluations, which limits their efficiency in applications.
@@
In contrast, one-step generators (Luo et al., 2024c, a; Zheng & Yang, 2024; Kang et al., 2023; Sauer et al., 2023a; Yin et al., 2024; Zhou et al., 2024a) have emerged as a highly efficient alternative to multi-step diffusion models. Unlike DMs, one-step generative models map latent noises directly to data samples in a single forward pass, making them highly efficient for real-time applications such as text-to-image and text-to-video generations on edge devices. Many existing works have demonstrated the leading performances of one-step text-to-image generators (Luo et al., 2024c; Zhou et al., 2024a; Yin et al., 2024) by employing diffusion distillation (Luo, 2023) as well as GAN techniques(Goodfellow et al., 2014; Brock et al., 2018; Karras et al., 2020; Sauer et al., 2023b).
  • Yet, these works only focus on matching the one-step model’s distributions with pre-trained diffusion models or ground truth data, overlooking the critical challenge of aligning one-step text-to-image models with human preferences, which is one of the central requirements of human-centric AI.
@@
Figure 1: Generated $1024\times 1024$ images from one-step 2.6B DI*-SDXL model. The DI*-SDXL-1step model outperforms FLUX-dev-50step(Lab, 2024) and SD3.5-large-28step(AI, 2024) in human preference metrics on the Parti prompt benchmark and HPSv2.1 scores. Please zoom in to check the details.
@@
To close this gap, we introduce Diff-Instruct* (DI*), a novel post-training approach that aligns one-step text-to-image generators with human preference while maintaining the ability to generate diverse and photo-realistic high-resolution images. We frame the human-preference alignment problem as reward maximization with score-based divergence constraints. This yields generated samples that not only have improved human reward but also adhere to user prompts. Unlike traditional RLHF methods (Christiano et al., 2017; Ouyang et al., 2022; Luo, 2024), which rely on Kullback-Leibler (KL) divergence for distribution regularization, we demonstrate that score-based regularization is important to preserve sample diversity and avoids reward hacking – the model generates weird, painting-like outputs with high rewards but low realism.
Our experiments demonstrate three key advantages: 1) Efficiency: The 2.6B DI*-SDXL-1step model generates $1024\times 1024$ resolution images in a single step, using only 1.88% of the inference time and 29.3% of the GPU memory of the 50-step 12B FLUX-dev model. 2) Superior Performance: On Parti prompts, DI*-SDXL-1step achieves higher PickScore, ImageReward, and CLIPScore than FLUX-dev-50step. It also sets a new state-of-the-art HPSv2.1 score of 31.19 among open-source models. 3) Fidelity: The model maintains diversity and prompt adherence, matching FLUX-dev and SD3.5-large on COCO benchmarks while being orders of magnitude faster.
We summarize our contributions as follows:
@@
  • We introduce Diff-Instruct* for preference alignment of one-step text-to-image generative models with a strict theoretical guarantee based on a new score-based PPO paradigm;
@@
  • We introduce two novel approaches to incorporate classifier-free guidance into human preference alignment through explicit-implicit reward decoupling.
@@
  • With extensive evaluations, our best open-sourced 2.6B DI*-SDXL-1step text-to-image model, a leading text-to-image model with a resolution of $1024\times 1024$ that significantly outperforms the leading FLUX-dev-50step model with 1.88% inference time.
@@ 2 Preliminary
Here, we introduce preliminary knowledge and notations about diffusion models. Let $q_{0}(\bm{x})=q_{d}(\bm{x})$ be the data distribution. The goal of generative modeling is to train models to generate new samples $\bm{x}\sim q_{0}(\bm{x})$. Under mild conditions, the forward diffusion process of a diffusion model can transform initial distribution $q_{0}$ towards some simple noise distribution,
@@
$\mathcal{L}_{DSM}(\varphi)$ $=\int_{t=0}^{T}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim q_{0},\atop\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\|\bm{s}_{\varphi}(\bm{x}_{t},t)$
$\quad-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$
The weighting function $\lambda(t)$ controls the importance of the learning at different time levels, and $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ denotes the conditional transition of the forward diffusion (2.1). After training, $\bm{s}_{\varphi}(\bm{x}_{t},t)\approx\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ is a good approximation of the marginal scores of the diffused data distribution.
3 Score-based Post-training
@@
Our basic setting is that we have a known human reward function $r(\bm{x}_{0},\bm{c})$, which encodes the human preference for an image $\bm{x}_{0}$ and corresponding text description $\bm{c}$. Besides, we also have a pre-trained diffusion model which will later act as a reference distribution $p_{ref}(\bm{x}_{0})=q_{0}(\bm{x}_{0})$. The reference diffusion model is specified by the score function $\bm{s}_{q_{t}}(\bm{x}_{t})\coloneqq\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$, where $q_{t}(\bm{x}_{t})$’s is the underlying distribution diffused at time $t$ according to (2.1). The pre-trained diffusion model represents the ground-truth data distribution that is used to prevent the one-step model from seeking high human reward but losing the ability to generate realistic images.
@@
Our goal is to train a human-preferred one-step text-to-image model $g_{\theta}$ that generates images by directly mapping a random noise $\bm{z}\sim p_{z}$ to obtain $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$, conditioned on the input text $\bm{c}\sim\mathcal{C}$. We want the one-step model’s output distribution, $p_{\theta}(\bm{x}_{0}|\bm{c})$, to maximize the expected human rewards while maintaining the ability to generate realistic images. We force $p_{\theta}(\bm{x}_{0}|\bm{c})$ to not diverge from $p_{ref}(\cdot)$ by using a divergence regularization term. Let $\bm{D}(\cdot,\cdot)$ be a distribution divergence. For any fixed prompt $\bm{c}$, the training objective is defined as:
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}|\bm{c})}\bigg{\{}\big{[}$ $-\alpha r(\bm{x}_{0},\bm{c})\big{]}+\bm{D}(p_{\theta},p_{ref})\bigg{\}}$ (3.1)
@@
Here $\alpha$ is a coefficient that balances reward influences and $\bm{D}(\cdot)$ acts as a regularization term. When using Kullback-Leibler divergence as the regularization term, the objective (3.1) turns to the online Proximal Policy Gradient (PPO) algorithm(Schulman et al., 2017), which has much success in reinforcement learning from human feedback (RLHF) for large language models (Ouyang et al., 2022) and one-step text-to-image models alignment (Luo, 2024). Particularly, recently (Luo, 2024) has shown that traditional KL-PPO can fairly train human-preferred one-step text-to-image models. However, KL divergence is notorious for mode-seeking behavior, which makes  Luo (2024)  very easy to collapse to some distribution modes – such as painting-like images – that have high human reward but lack diversity. To address this issue, we find that general score-based divergences have an appealing diversity-preserving property that has been proved in one-step diffusion distillation (Luo et al., 2024c; Zhou et al., 2024b). Based on this intuition, we propose Diff-Instruct*, a novel post-training approach built upon a new Score-based online PPO paradigm that uses a general score-based divergence to regularize the RLHF process.
3.1 Score-based Divergences
@@
Inspired by Score-Implicit Matching (SIM (Luo et al., 2024c)) theory that revealed the advantages of general score-based divergences over the traditional KL divergence, we define the regularization term $\bm{D}(p_{\theta},p_{ref})$ in (3.1) via the following general score-based divergence. Assume $\mathbf{d}:\mathbb{R}^{d}\to\mathbb{R}$ is a scalar-valued proper distance function (i.e., a non-negative function that satisfies $\forall\bm{x},\mathbf{d}(\bm{x})\geq 0$ and $\mathbf{d}(\bm{x})=0$ if and only if $\bm{x}=\bm{0}$). Given a parameter-independent sampling distribution $\pi_{t}$ that has large distribution support, we can formally define a time-integral score divergence as
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})\coloneqq\int_{0}^{T}w(t)\mathbb{E}_{\pi_{t}}\mathbf{d}\big{(}\bm{s}_{p_{\theta,t}}-\bm{s}_{q_{t}}\big{)}\mathrm{d}t.$ (3.2)
@@
3.2 Theories for Tractable Objectives
Recall that $g_{\theta}$ is a one-step model, therefore, samples from $p_{\theta}$ can be implemented through a direct and differentiable mapping $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$. With the score-based regularization term (3.2), for each given text prompt $\bm{c}$, we can formally write down our training objective to minimize as:
@@ 3.2 Theories for Tractable Objectives
$\quad+\int_{t=0}^{T}w(t)\mathbb{E}_{\bm{x}_{t}|\bm{x}_{0}\atop\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\bigg{\{}-\mathbf{d}^{\prime}(\bm{y}_{t})\bigg{\}}^{T}$
$\quad\bigg{\{}\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}\mathrm{d}t\bigg{]}.$
@@
Under mild assumptions, if we take the sampling distribution in (3.2) as $\pi_{t}=p_{\operatorname{sg}[\theta],t}$, then the gradient of (3.3) w.r.t $\theta$ is the same as (3.4):
$\frac{\partial}{\partial\theta}\mathcal{L}_{Orig}(\theta)=\frac{\partial}{\partial\theta}\mathcal{L}_{DI*}(\theta).$
In practice, we can use another assistant diffusion model $\bm{s}_{\psi}(\bm{x}_{t},t)$ to approximate the one-step model’s score function $\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})$ pointwise, which was also done in the literature of diffusion distillations works such as Luo et al. (2024b, c); Zhou et al. (2024a, b); Yin et al. (2023, 2024). Therefore, we can alternate between 1) updating the assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t},t)$ using one-step model generated samples (which are efficient) and 2) updating the one-step model by minimizing the tractable objective (3.4). We name our training method that minimizes the objective $\mathcal{L}_{DI*}(\theta)$ in (3.4) the Diff-Instruct* because it is inspired by Diff-Instruct (Luo et al., 2024b) and Diff-Instruct++ (Luo, 2024) that involves an additional diffusion model and a reward model to train one-step text-to-image models.
3.3 Understanding Classifier-free Guidance
Input: prompt dataset $\mathcal{C}$, generator $g_{\theta}(\bm{x}_{0}|\bm{z},\bm{c})$, prior distribution $p_{z}$, reward model $r(\bm{x},\bm{c})$, reward model scale $\alpha_{rew}$, CFG reward scale $\alpha_{cfg}$, reference diffusion model $\bm{s}_{ref}(\bm{x}_{t}|c,\bm{c})$, assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})$, forward diffusion $p(\bm{x}_{t}|\bm{x}_{0})$ (2.1), assistant diffusion updates rounds $K_{TA}$, time distribution $\pi(t)$, diffusion model weighting $\lambda(t)$, generator IKL loss weighting $w(t)$.